OpenStudy (anonymous):

Circle A has an equation of x2 + (y + 3)2 = 16. Where is the center of the circle located? on the x-axis in the third quadrant on the y-axis in the first quadrant

OpenStudy (anonymous):

anyone please assist me?

OpenStudy (accessdenied):

The center of a circle in the form \((x - h)^2 + (y - k)^2 = r^2\) is the point \((h, \; k)\). \(x^2 + (y + 3)^2 = 16\) The y + 3 = y - (-3), so we can see that the y-value is -3. (x, -3) While there isn't anything with the x, consider replacing x by x+0. That may help. :)

OpenStudy (anonymous):

C?

OpenStudy (anonymous):

i mean on the x axis?

OpenStudy (accessdenied):

Would you agree that the point is (0, -3)? The x-value is 0, so does not move in the direction of the x-axis. It has y-value -3, so it moves down from the origin 3 units. That would be the y-axis, correct? I think you had it right the first time.

OpenStudy (anonymous):

yeah the y axis sorry thanks so much

OpenStudy (accessdenied):

You're welcome. :)